Circular No. 8214

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003im, 2003in, 2003io, 2003ip Further to IAUC 8207, W. Li reports the LOTOSS discovery of three

apparent supernovae (2003im, 2003in, 2003ip) on unfiltered KAIT images. Also, T. Puckett, Mountain Town, GA; and B. Kerns, Stanton, ND, report the discovery of an apparent supernova (2003io) on an unfiltered CCD frame (limiting mag 19.3) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope, the new object being confirmed on unfiltered CCD frames taken by M. Marcus on Oct. 2.16 UT.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003im Sept. 30.4 0^{h}44^{m}59.^{s}28 − 8^{o}53′11′′.5 17.4 1′′.6 E, 11′′.6 N
2003in Sept. 30.5 3 35 33.31 + 5 03 52.7 17.8 2′′.9 E, 8′′.5 S
2003io Sept. 30.27 2 17 26.87 +14 34 36.3 17.1 11′′.3 E, 12′′.5 S
2003ip Oct. 1.3 0 33 16.76 + 7 54 19.5 16.5 20′′.3 E, 1′′.2 N

Puckett adds that SN 2003io in UGC 1761 does not appear on unfiltered images taken by himself on 1999 Oct. 17 or 2000 Jan. 2 and 7 (limiting mag ∼ 20.0). Additional approximate LOTOSS magnitudes (K = KAIT image; T = Tenagra II image): SN 2003im in a galaxy within Abell cluster 85, 2002 Dec. 28.1, [19.0 (T); Oct. 1.4, 17.3 (K). SN 2003in in IC 1956, Jan. 23.1, [19.5 (T); Oct. 1.5, 17.7 (K). SN 2003ip in UGC 327, 2002 Dec. 29.1, [19 (T); 2003 Oct. 2.3, 16.5 (K).

A. V. Filippenko and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Oct. 2 UT with the Shane 3-m telescope at Lick Observatory, reveal that SN 2003im and SN 2003in are of type Ia, roughly 1 month and 2–3 weeks past maximum brightness, respectively. SN 2003ip is a supernova of type II, within a few weeks after the explosion; hydrogen Balmer lines with P-Cyg profiles are superposed on a relatively blue continuum, but the emission component of Hα substantially dominates over the absorption component. (The offset above for SN 2003in is a correction by Filippenko to that given on CBET 45.)

SUPERNOVAE 2003ij AND 2003ik Filippenko and Foley add that inspection of CCD spectra, obtained as

above on Oct. 2 UT, reveal that SN 2003ij (IAUC 8210) and SN 2003ik (IAUC 8212) are also of type Ia, within a few days of maximum brightness.

2003 October 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green